OpenStudy (kayla2001):

what is 12ft and 6inches divided by 928ft and 2 inches

OpenStudy (samanthagreer):

((12 ft) + (6 inches)) / ((928 ft) + (2 inches)) = 0.0134674089

OpenStudy (samanthagreer):

did that help

OpenStudy (kayla2001):

I got another question

OpenStudy (kayla2001):

my house is 928 ft and 2 inches the size of the siding is 12 ft and 2 inche how many pieces do I need?

OpenStudy (phi):

do you mean siding is 12 ft, 6 inches ?

OpenStudy (kayla2001):

yes

OpenStudy (phi):

I would change 6" to feet. do you know how to do that ?

OpenStudy (kayla2001):

no

OpenStudy (phi):

there are 12 inches in one foot

OpenStudy (phi):

write \[ \frac{1 \ foot}{12\ inches} \] and multiply by 6 inches: \[6\ inches\cdot \frac{1 \ foot}{12\ inches} \] and simplify

OpenStudy (kayla2001):

my answer I got was 72